OpenStudy (anonymous):

1. A manufacturing company manufactures a cardboard box with a square base and a height of 15 inches. Suppose the equation x^2+60x-7200=0 can be used to find the length and width of the base of the box, each measuring x inches.

OpenStudy (anonymous):

-write the quadratic equation in factored form -use the zero product property to solve the equation -explain how the solution relates to the situation

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

factored form (x-60)(x+120)=0

rebeccaxhawaii (rebeccaxhawaii):

okay so what have you got so far?

OpenStudy (anonymous):

-use the zero product property to solve the equation (x-60)(x+120)=0 x-60=0 x+120=0 x-60+60 = 0+60x+120-120 = 0-120 x=60 x=-120

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

@4everaddicted2anime

OpenStudy (4everaddicted2anime):

Yes that is correct

OpenStudy (anonymous):

-explain how the solution relates to the situation

OpenStudy (4everaddicted2anime):

Well the manufacturing company needs to know how much the box can hold(area) to price it right?

OpenStudy (anonymous):

so, the solution = the area of the boxes which will help the manufactouring company price there products at a resenible prise.

OpenStudy (4everaddicted2anime):

yes

OpenStudy (anonymous):

so thatll work for an awser

OpenStudy (4everaddicted2anime):

yeah, just put volume instead of area(I miss-typed when I said area earlier)

OpenStudy (anonymous):

ok, thanks

OpenStudy (4everaddicted2anime):

yw